OpenStudy (openstudyoperator):

The fisherman’s ropes are proportional in their length and weight. A rope the length of 30 m has a weight of 2 kg. Use this proportion of length to weight to determine the weight of a rope with the length of 105 m. The 105 m rope weighs ______kg. Question 14 options: A) 52.5 kg B) 7 kg C) 28 kg D) 15 kg

OpenStudy (ethanbradberry):

should be 7 and ill tell you how i got that, devide 2 by 30 to get the ratio which is 0.66666 and then multiply 0.66666 by 105 and you get 7

OpenStudy (ethanbradberry):

if this helped plz medal and fan :)

OpenStudy (irishboy123):

\(l \propto W\) \(\implies l = k W\) \(\implies \dfrac{l_1}{l_2} = \dfrac{ W_1}{W_2}\) \(\implies \dfrac{30}{105} = \dfrac{ 2}{W_2}\) \(\implies W_2 = \dfrac{2 \cdot 105}{30} = 7\)